# Bright ring features and polarization structures in Kerr-Sen black hole images of Sagittarius A* illuminated by radiatively inefficient accretion flows

### v3

@@ III EXTRACTED RING FEATURES in images of Kerr-Sen black holes
To compare theoretical simulations with the photos captured by the EHT under the current observational precision, it is necessary and critical to extract bright ring features from blurred black hole images generated by GRRT. In the REX algorithm employed by the EHT EventHorizonTelescope:2019ths, the ring center $(x_{0},y_{0})$ is identified by the position that minimizes the normalized radial peak dispersion
@@
With this REX algorithm, we present visualization of extracted ring features of blurred images for different dilaton parameters $r_{2}$ and disk thickness $H$ in Fig. 2. Moreover, we also present the corresponding unwrapped ring profiles of the simulated images of Kerr-Sen black holes in 3. Our results show that the diameter of the bright ring decreases with the dilaton parameter $r_{2}$, which is similar to that in the unblurred case shown in Fig. 1. However, the width and brightness of the ring increase with the dilaton parameter $r_{2}$. With the increase of the disk thickness $H$, we find that the diameter and width of the bright ring decrease. The brightness also increases with disk thickness $H$, but the brightness enhancement induced by the disk thickness is less prominent than that driven by the dilaton parameter $r_{2}$.
From the observation of EHT sgraI; sgraIV; EHTsgraVI; 2022ApJ...930L..16E; Chael_2018, the bright ring diameter of the supermassive black hole Sgr A* is $d_{\rm ring}=51.8\pm 2.3~\mu{\rm as}$. Taking into account the current precision of observation, one can compare the bright ring diameter obtained from the EHT observations with the bright ring diameter extracted from the simulated blurred images of Kerr-Sen black holes under a Gaussian beam of FWHM 2020 $\mu$.
@@
In Fig. 4, we present distribution of the bright ring size on the $r_{2}-a$ plane for the blurred images of Kerr-Sen black holes with different inclination $i$ and disk thicknesses $H$. The gray triangles represent results of the ring diameters extracted by the REx algorithm from the GRRT simulations. The solid and dashed black lines respectively correspond to $d=51.8\mu as$ and its $1\sigma$ uncertainty range (i.e., $\pm 2.3\mu as$), which are from the observation of EHT. The blue contours correspond to regions with black hole parameters where the angular momentum at the outer event horizon exceeds that of Kerr black holes with $a=0.5$, i.e., $\Omega_{H}>0.134$.
Figure 5: Distribution of the bright ring size on the $r_{2}-a$ plane for the unblurred images of Kerr-Sen black holes with different inclination $i$ and disk thicknesses $H$. Others are same as in Fig. 3.
@@
Figure 6: Comparison of the horizontal cross-sectional normalized flux density between images of Kerr black holes and Kerr-Sen black holes with the same ring diameter $d_{\rm ring}$ for different disk thickness at fixed inclination angle $i=30^{\circ}$.
@@
To gain a more detailed understanding of the image of Sgr A* in the context of the Kerr-Sen black hole, we further explored, in Figs. 6 and 7, the differences in normalized flux density between the images of Kerr black holes and Kerr-Sen black holes under different thicknesses $H$ and inclinations $i$, with each image having the same ring diameter $d_{\rm ring}=51.8~\mu{\rm as}$. Figs. 6 and 7 represents the normalized flux density along the horizontal cross section of the images. As the disk thickness increases, the difference in profiles between the images of Kerr black holes and Kerr-Sen black holes decreases, and the profiles almost coincide when $H=0.3$. This means that for black hole images with the same ring diameter $d_{\rm ring}$,it is more difficult to distinguish between Kerr black hole and Kerr-Sen black hole based on the brightness distribution when the accretion disk has a larger thickness. In addition, from Fig. 7, we find that the difference in profiles decreases with the inclination angle $i$, such that the profiles almost coincide when $i=30^{\circ}$. Comparing with Figs. 6 and 7, we can find that effect of the accretion disk thickness on the difference in profiles is slightly more pronounced than that of the inclination angle. Therefore, to observe these differences and further resolve the degeneracy in images, higher-resolution observations are required, such as those that will be obtainable from future ngEHT observations tiede2022measuringphotonringsngeht; Johnson_2023.
@@ IV Polarization structures in images of Kerr-Sen black holes
In general, the polarization information carried by light beams can be described by the Stokes parameters $I$, $Q$, $U$, and $V$ 1996AAS..117..161H; Smirnov_2011. The parameter $I$ represents the total emission, and other parameters $Q$, $U$, and $V$ describe the polarized portion of the emission. Using these quantities, the electric-vector position angle (EVPA) is defined as
@@
To further quantitatively characterize the polarization mode distribution, one can define a set of quantitative indices $\beta_{m}$ for the EVPA patterns through the scalar product of the complex polarization quantity $P=Q+iU$ and a set of basis functions ehtVII; m87VIII; m87VIII; Palumbo_2020; Chael_2023, i.e.,
$\beta_{m}=\frac{1}{I_{\rm ann}}\int_{r_{min}}^{r_{max}}\int_{0}^{2\pi}P(r,\phi)e^{im\phi}r\mathrm{~d}\phi\mathrm{~d}r$, (14)
where the chosen basis functions $P_{b}=e^{im\phi}$ represent the possible spatial modes of EVPA variation with the azimuthal angle $\phi$ in the image plane, and $m$ is the corresponding mode order m87VIII; Palumbo_2020; Chael_2023. The quantity $I_{\rm{ann}}$ has a form
$I_{\rm{ann}}=\int_{r_{min}}^{r_{max}}\int_{0}^{2\pi}I(r,\phi)r\mathrm{d}\phi\mathrm{d}r.$ (15)
where the inner radius $r_{min}$ and outer radius $r_{max}$ correspond to the inner and outer edges of the emission ring shown in the image, respectively. In general, the $\beta_{m}$ coefficients are complex numbers. The magnitude $|\beta_{m}|$ quantifies the degree of $m$-fold rotational symmetry in the EVPA distribution around the ring, while the argument $\angle\beta_{m}$ measures the average orientation of the symmetric component, relative to a reference pattern with vertical EVPA at image north m87VIII; Palumbo_2020. The model $m=2$ serves as a robust discriminator for distinguishing between MAD and SANE flows in terms of black hole images.
@@
Figs. 9 and 10 present changes of the complex mode $\beta_{2}$ with the normalized dilaton parameter $r_{2}/r_{2\text{max}}$ for fixed inclination $i=30^{\circ}$ and different disk thickness $H$ after applying a 20 $\mu as$ Gaussian blur to the original simulated images of SgrA*. As the dilaton parameter increases, the real part $Re\beta_{2}$ increases and the imaginary part $Im\beta_{2}$ decreases. Moreover, it is found that the real part $Re\beta_{2}$ is always negative, which means the phase angle is in the range $\angle\beta_{2}\in(\pi/2,\pi)$. The magnitude $|\beta_{2}|$ generally exhibits a declining trend as $r_{2}/r_{2\text{max}}$ increases. Particularly in the case of a low spin parameter, the decline is more pronounced. In contrast, in the case with large spin parameter, the decline is slower or even tends to flatten. The angle $\angle\beta_{2}$ increases monotonically with $r_{2}/r_{2\text{max}}$, while its rate of increase decreases with the black hole spin. As the disk thickness $H$ increases, the magnitude $|\beta_{2}|$ decreases slightly, and the phase angle $\angle\beta_{2}$ increases slightly. However, these effects are much weaker than those induced by the dilaton parameter. These results are beneficial in understanding black hole images with accretion disks and the gravity of Einstein-Maxell-Dailton-Axion theory.
@@ V Summary
We have investigated the bright ring features and polarization structures in the images illuminated by the 230 GHz thermal synchrotron emission from the RIAF around Kerr-Sen black holes. Our results show that as the dilaton parameter increases, bright ring shrinks, while the width and brightness of the ring increase. With the increase of the disk thickness, the diameter and width of the bright ring decrease. The brightness also increases with disk thickness $H$, but the brightness enhancement induced by the disk thickness is less prominent than that driven by the dilaton parameter $r_{2}$.. We also present the allowed ranges of black hole parameters through comparing with the EHT observational data. For blurred images convolved with a Gaussian kernel $20\,\mu\mathrm{as}$, we find that for a fixed disk thickness $H$, increasing the inclination angle $i$ reduces the lower bounds of parameters $r_{2}$ and $a$, while their upper bounds remain largely unchanged. For a fixed inclination $i$, increasing the disk thickness $H$ leads to decreasing the lower bounds of $r_{2}$ and $a$, and shifting the overall allowed parameter space toward the region with the lower values of $a$ and $r_{2}$. These imply that the disk thickness $H$ exerts a stronger influence on the parameter space than the observer’s inclination $i$, and a larger disk thickness $H$ tends to yield a smaller bright ring diameter. In contrast, for the unblurred images, we find that there exhibits the higher bottom bounds for $a$ and $r_{2}$, while the upper bound of $a$ slows down and the upper bound of $r_{2}$ increases. These also confirm that the observation resolution affects significantly the precision of constraining black hole parameters.
@@
We also analyzed the influence of the dilaton parameter $r_{2}$ on polarization structures in Kerr-Sen black hole images by quantifying EVPA rotational symmetry using the method of $\beta_{2}$. Our results show that as $r_{2}$ increases, the real part $Re\beta_{2}$ increases and the imaginary part $Im\beta_{2}$ decreases, but the magnitude $|\beta_{2}|$ generally exhibits a declining trend. The angle phase $\angle\beta_{2}$ increases monotonically and its increasing rate decreases with the black hole spin. The disk thickness $H$ results in that both the magnitude $|\beta_{2}|$ decreases slightly, and the phase angle $\angle\beta_{2}$ increases slightly, but its effects are much weaker than those from the dilaton parameter.

### v4

@@ III EXTRACTED RING FEATURES in images of Kerr-Sen black holes
To compare theoretical simulations with the photos captured by the EHT under the current observational precision, it is necessary and critical to extract bright ring features from blurred black hole images generated by GRRT. In the REX algorithm employed by the EHT Akiyama and others (2019a), the ring center $(x_{0},y_{0})$ is identified by the position that minimizes the normalized radial peak dispersion
@@
With this REX algorithm, we present the visualization of extracted ring features of blurred images for different dilaton parameters $r_{2}$ and disk thickness $H$ in Fig. 2. Moreover, we also present the corresponding unwrapped ring profiles of the simulated images of Kerr-Sen black holes in Fig. 3. Our results show that the diameter of the bright ring decreases with the dilaton parameter $r_{2}$, which is similar to that in the unblurred case shown in Fig. 1. However, the width and brightness of the ring increase with the dilaton parameter $r_{2}$. With the increase of the disk thickness $H$, we find that the diameter and width of the bright ring decrease. The brightness also increases with disk thickness $H$, but the brightness enhancement induced by the disk thickness is less prominent than that driven by the dilaton parameter $r_{2}$.
From the observation of EHT Collaboration et al. (2022b, c, a); Event Horizon Telescope Collaboration et al. (2022); Chael et al. (2018), the bright ring diameter of the supermassive black hole Sgr A* is $d_{\rm ring}=51.8\pm 2.3~\mu{\rm as}$. Taking into account the current precision of observation, one can compare the bright ring diameter obtained from the EHT observations with the bright ring diameter extracted from the simulated blurred images of Kerr-Sen black holes under a Gaussian beam of FWHM 2020 $\mu$.
@@
In Fig. 4, we present distribution of the bright ring size on the $r_{2}-a$ plane for the blurred images of Kerr-Sen black holes with different inclination $i$ and disk thicknesses $H$. The gray triangles represent results of the ring diameters extracted by the REx algorithm from the GRRT simulations. The solid and dashed black lines respectively correspond to $d=51.8\mu as$ and its $1\sigma$ uncertainty range (i.e., $\pm 2.3\mu as$), which are from the observation of EHT. The blue contours denote regions where outer event horizon angular momentum satisfies $\Omega_{\rm H}\geq 0.134$. The lower limit of $0.134$ corresponds to the outer event horizon angular momentum of Kerr black holes with $a=0.5$, given that previous analyses of Sgr A* images favor a relatively high black hole spins Event Horizon Telescope Collaboration et al. (2022).
Figure 5: Distribution of the bright ring size on the $r_{2}-a$ plane for the unblurred images of Kerr-Sen black holes. The parameter settings for the inclination $i$ (rows) and disk thickness $H$ (columns) are the same as in Fig. 4.
@@
Figure 6: Comparison of the horizontal cross-sectional normalized flux density between images of Kerr black holes and Kerr-Sen black holes with the same ring diameter $d_{\rm ring}$ for different disk thickness at fixed inclination $i=30^{\circ}$.
@@
To gain a more detailed understanding of the image of Sgr A* in the context of the Kerr-Sen black hole, we further explored, in Figs. 6 and 7, the differences in normalized flux density between the images of Kerr black holes and Kerr-Sen black holes under different thicknesses $H$ and inclinations $i$, with each image having the same ring diameter $d_{\rm ring}=51.8~\mu{\rm as}$. Figs. 6 and 7 represents the normalized flux density along the horizontal cross section of the images. As the disk thickness increases, the difference in profiles between the images of Kerr black holes and Kerr-Sen black holes decreases, and the profiles almost coincide when $H=0.3$. This means that for black hole images with the same ring diameter $d_{\rm ring}$,it is more difficult to distinguish between Kerr black hole and Kerr-Sen black hole based on the brightness distribution when the accretion disk has a larger thickness. In addition, from Fig. 7, we find that the difference in profiles decreases with the inclination angle $i$, such that the profiles almost coincide when $i=30^{\circ}$. Comparing with Figs. 6 and 7, we can find that effect of the accretion disk thickness on the difference in profiles is slightly more pronounced than that of the inclination angle. Therefore, to observe these differences and further resolve the degeneracy in images, higher-resolution observations are required, such as those that will be obtainable from future ngEHT observations Tiede et al. (2022); Johnson et al. (2023).
@@ IV Polarization structures in images of Kerr-Sen black holes
In general, the polarization information carried by light beams can be described by the Stokes parameters $I$, $Q$, $U$, and $V$ Hamaker and Bregman (1996); Smirnov (2011). The parameter $I$ represents the total emission, and other parameters $Q$, $U$, and $V$ describe the polarized portion of the emission. Using these quantities, the electric-vector position angle (EVPA) is defined as
@@
To further quantitatively characterize the polarization mode distribution, one can define a set of quantitative indices $\beta_{m}$ for the EVPA patterns through the scalar product of the complex polarization quantity $P=Q+iU$ and a set of basis functions Akiyama et al. (2021a, b, b); Palumbo et al. (2020); Chael et al. (2023), i.e.,
$\beta_{m}=\frac{1}{I_{\rm ann}}\int_{r_{\rm min}}^{r_{\rm max}}\int_{0}^{2\pi}P(r,\phi)e^{im\phi}r\mathrm{~d}\phi\mathrm{~d}r,$ (15)
where the chosen basis functions $P_{b}=e^{im\phi}$ represent the possible spatial modes of EVPA variation with the azimuthal angle $\phi$ in the image plane, and $m$ is the corresponding mode order Akiyama et al. (2021b); Palumbo et al. (2020); Chael et al. (2023). The quantity $I_{\rm{ann}}$ has a form
$I_{\rm{ann}}=\int_{r_{\rm min}}^{r_{\rm max}}\int_{0}^{2\pi}I(r,\phi)r\mathrm{d}\phi\mathrm{d}r.$ (16)
where the inner radius $r_{\rm min}$ and outer radius $r_{\rm max}$ correspond to the inner and outer edges of the emission ring shown in the image, respectively. In general, the $\beta_{m}$ coefficients are complex numbers. The magnitude $|\beta_{m}|$ quantifies the degree of $m$-fold rotational symmetry in the EVPA distribution around the ring, while the argument $\angle\beta_{m}$ measures the average orientation of the symmetric component, relative to a reference pattern with vertical EVPA at image north Akiyama et al. (2021b); Palumbo et al. (2020). The model $m=2$ serves as a robust discriminator for distinguishing between MAD and SANE flows in terms of black hole images.
@@
Figs. 9 and 10 present changes of the complex mode $\beta_{2}$ with the normalized dilaton parameter $r_{2}/r_{2\text{max}}$ for fixed inclination $i=30^{\circ}$ and different disk thickness $H$ after applying a 20 $\mu as$ Gaussian blur to the original simulated images of SgrA*. As the dilaton parameter increases, the real part $Re\beta_{2}$ increases and the imaginary part $Im\beta_{2}$ decreases. Moreover, it is found that the real part $Re\beta_{2}$ is always negative, which means the phase angle is in the range $\angle\beta_{2}\in(\pi/2,\pi)$. The magnitude $|\beta_{2}|$ generally exhibits a declining trend as $r_{2}/r_{2\text{max}}$ increases. Particularly in the case of a low spin parameter, the decline is more pronounced. In contrast, in the case with large spin parameter, the decline is slower or even tends to flatten. The angle $\angle\beta_{2}$ increases monotonically with $r_{2}/r_{2\text{max}}$, while its rate of increase decreases with the black hole spin. As the disk thickness $H$ increases, the magnitude $|\beta_{2}|$ decreases slightly, and the phase angle $\angle\beta_{2}$ increases slightly. However, these effects are much weaker than those induced by the dilaton parameter. These results are beneficial in understanding black hole images with accretion disks and the gravity of Einstein-Maxell-Dilaton-Axion theory.
@@ V Summary
We have investigated the bright ring features and polarization structures in the images illuminated by the 230 GHz thermal synchrotron emission from the RIAF around Kerr-Sen black holes. Our results reveal that an increase in the dilaton parameter leads to a shrinking of the bright ring, accompanied by enhancements in both its width and brightness. Actually, these effects can be explained by the fact that the dilaton parameter $r_{2}$ weakens the black hole’s gravity because the outer horizon radius $r_{+}$ decreases with the dilaton parameter $r_{2}$ due to $\frac{dr_{+}}{dr_{2}}=-\frac{r_{+}}{2\sqrt{(M-\frac{r_{2}}{2})^{2}-a^{2}}}<0$. Therefore, compared to black holes without dilaton parameter $r_{2}$, only photons passing closer to the black hole can fall into it, thereby reducing the diameter of the bright ring. Simultaneously, it allows light emitted from broader regions on the accretion disk to reach the observer, leading to an increase in both the brightness and width of the bright ring. In parallel, we find that as the disk thickness grows, both the diameter and width of the bright ring decrease. While the brightness also increases with the disk thickness $H$, this enhancement is less prominent than that driven by the dilaton parameter $r_{2}$. We also present the allowed ranges of black hole parameters through comparing with the EHT observational data. For blurred images convolved with a Gaussian kernel $20\,\mu\mathrm{as}$, we find that for a fixed disk thickness $H$, increasing the inclination angle $i$ reduces the lower bounds of parameters $r_{2}$ and $a$, while their upper bounds remain largely unchanged. For a fixed inclination $i$, increasing the disk thickness $H$ leads to decreasing the lower bounds of $r_{2}$ and $a$, and shifting the overall allowed parameter space toward the region with the lower values of $a$ and $r_{2}$. These imply that the disk thickness $H$ exerts a stronger influence on the parameter space than the observer’s inclination $i$, and a larger disk thickness $H$ tends to yield a smaller bright ring diameter. In contrast, for the unblurred images, we find that there exhibits the higher bottom bounds for $a$ and $r_{2}$, while the upper bound of $a$ slows down and the upper bound of $r_{2}$ increases. These also confirm that the observation resolution affects significantly the precision of constraining black hole parameters.
@@
We also analyzed the influence of the dilaton parameter $r_{2}$ on polarization structures in Kerr-Sen black hole images by quantifying EVPA rotational symmetry using the method of $\beta_{2}$. Our results show that as $r_{2}$ increases, the real part $Re\beta_{2}$ increases and the imaginary part $Im\beta_{2}$ decreases, but the magnitude $|\beta_{2}|$ generally exhibits a declining trend. The angle phase $\angle\beta_{2}$ increases monotonically and its increasing rate decreases with the black hole spin. In refs.Wong et al. (2025); Chael (2025), one can find that the phase angle $\angle\beta_{2}$ from the 230 GHz images from the radiative simulations can be approximately fitted with a simple function $\angle\beta_{2}=-2\arctan\frac{C_{0}}{|\Omega_{H}|}+C_{1}$, where $\Omega_{H}=\frac{a}{2r_{+}}$ is the angular velocity at outer event horizon, $C_{0}$ and $C_{1}$ are numerical parameters. Consequently, one has the value $|\beta_{2}|\sim\sqrt{1+\tan^{2}\angle\beta_{2}}\sim\frac{2\Omega_{H}}{\Omega^{2}_{H}-C^{2}_{0}}$. With these fitted formulas, one can find that an increase in angular velocity triggered by the dilaton parameter results in a corresponding increase in the angle $\angle\beta_{2}$ and a decrease in the magnitude $|\beta_{2}|$. This further highlights the importance of the frame-dragging effect with respect to the polarization structure in black hole images. The disk thickness $H$ results in that both the magnitude $|\beta_{2}|$ decreases slightly, and the phase angle $\angle\beta_{2}$ increases slightly, but its effects are much weaker than those from the dilaton parameter.